OpenStudy (anonymous):

a A person deposited $500 in a savings account that pays 5% annual interest that is compounded yearly. At the end of 10 years, how much money will be in the savings account?

Directrix (directrix):

A = P ( 1 + r/n)^(nt) where A is amount, P is principal, r = rate, n= number of compounding periods per year and t = number of years. A = 500(1 + .05/1) ^ (1*10) A = $ 814.45. I checked my answer here at the Online Calculator. Check out the online compound interest calculator at http://www.webmath.com/compinterest.html